Q(P)=2392-26P

Calculate demand when P=$45

Calculate the change in quantity demanded if the price increases by $1

Calculate the price where the quantity demanded will equal zero
Answer #1

Answer

Q(P)=2392-26P

a) When P = 45, Q(45)=2392-26*45 = 1222

Here P represents Price and Q represents demand.

Hence when Price(P) = 45, Demand (Q) = 1222

b)

Q =2392-26P => \Delta Q=\Delta (2392-26P) = \Delta 2392 - \Delta 26P = -26\Delta P

Here Price increases by $1 => \Delta Q = -26*1 = -26 (negative sign suggest there quantity demanded decreases)

Hence If Price increases by $1, quantity demanded will decrease by 26

c)

Suppose Quantity demanded = 0 => Q(P) =Q =0

Hence Q(P) = 2392-26P = 0

=> P = 2392/26 = $92

Hence, At Price = $92, quantity demanded will equal to zero, Hence answer is Price(P) = $92